# TinyGPT-V: Efficient Multimodal Large Language Model via Small Backbones

### v1

@@ 4 Experiments
Figure 4: Changes in learning rate during the training stage of TinyGPT-V.
@@ 4.1 Training
Training process. In our experimental process, we meticulously orchestrated the training of our model through four distinct stages, each characterized by specific learning rate strategies and loss profiles, as shown in Figure 4 and Figure 5.
Stage 1: Spanning 17 epochs, with each epoch consisting of 1000 iterations, we employed a dynamic learning rate approach. The learning rate commenced at 1e-5 at the beginning of each epoch and gradually ascended to 1e-4 by the epoch’s end. This pattern was consistently applied across all 17 epochs. The training loss exhibited a steady decline, starting from 7.152 and progressively tapering down to 2.620, reflecting the model’s increasing proficiency in learning from the data. The purpose of this stage is to be able to make the Phi-2 model in TinyGPT-V react in some way to the input of the image modality. The alignment of text and image in the semantic space is done.
Stage 2: Comprising 4 epochs, each with 5000 iterations, this stage introduced the “linear_warmup_cosine_lr“ he2018bag ; goyal2018accurate learning rate schedule. We initiated a warmup phase of 5000 steps, where the learning rate linearly increased from 1e-6 (warmup_lr) to 1e-4 (init_lr), followed by a cosine decay down to a minimum learning rate of 8e-5. This phase saw a consistent reduction in loss, starting at 2.726 and culminating at 2.343. The purpose of this stage is to enable the LoRA module to play a role in multimodal data, further reducing the model’s loss on image-text pairs and improving the model’s ability to learn from the data.
@@
Table 2: Comparative performance of TinyGPT-V and other MLLMs across multiple visual question answering benchmarks.
Method Parameters Grounding GQA VSR IconVQ VizWiz HM
(zero-shot) (zero-shot) (zero-shot) (zero-shot)
Flamingo 9B - 31.8 - 28.8 57.0
BLIP-2 13B 41.0 50.9 40.6 19.6 53.7
LLaVA 13B 41.3 51.2 43.0 - -
Shikra 13B - - - - -
InstructBLIP 13B 49.5 52.1 44.8 33.4 57.5
MiniGPT-4 13B 30.8 41.6 37.6 - -
@@
TinyGPT-V 2.8B 33.6 53.2 43.3 24.8 53.2
@@ 4.2 Evaluation
Evaluation datasets. GQA hudson2019gqa is a dataset for real-world visual reasoning and compositional question answering, featuring a powerful question engine that generates 22 million diverse reasoning questions. VSR liu2023visual comprises over 10k natural text-image pairs in English, encompassing 66 types of spatial relations. IconQA lu2022iconqa with 107,439 questions aimed at challenging visual understanding and reasoning in the context of icon images, encompassing three sub-tasks (multi-image-choice, multi-text-choice, and filling-in-the-blank). VizWiz gurari2018vizwiz is a collection of more than 31,000 visual queries, each derived from a photo taken by a visually impaired individual using a smartphone, accompanied by a vocalized question regarding the image, and supplemented with 10 answers sourced from a crowd for each query. The Hateful Memes dataset (HM) kiela2021hateful , developed by Facebook AI, is a comprehensive multimodal collection specifically designed for the detection of hateful content in memes, combining both image and text elements, and comprises over 10,000 newly created multimodal examples.
@@
Visual question answering results. As shown in Table 2, it becomes evident that TinyGPT-V, a model with only 2.8 billion parameters, exhibits notably competitive performance across multiple benchmarks, closely rivaling models with nearly 13 billion parameters. Specifically, in the VSR (Visual Spatial Reasoning) zero-shot task, TinyGPT-V outshines its counterparts by securing the highest score of 53.2%. This is particularly impressive considering its parameter size is approximately 4.6 times smaller than other leading models such as BLIP-2, LLaVA, and InstructBLIP. In the GQA benchmark, while TinyGPT-V scores 33.6%, it lags behind the highest score achieved by InstructBLIP, which is 49.5%. However, TinyGPT-V shows robust performance in the IconVQ challenge, attaining a score of 43.3%, just 1.5% short of InstructBLIP’s leading score of 44.8%. Similarly, in the VizWiz task, TinyGPT-V demonstrates commendable capabilities with a score of 24.8%, which, though not the highest, is notable given its reduced parameter count. In the context of the Hateful Memes (HM) dataset, TinyGPT-V matches InstructBLIP’s top score of 57.5% with its own score of 53.2%, again underscoring its efficiency and capacity to compete with models of larger scales. Overall, TinyGPT-V’s performance across these diverse and challenging benchmarks is striking, especially when considering its parameter efficiency
@@ 4.3 Ablation Study
Method Stage 1 Loss Stage 2 Loss Stage 3 Loss Stage 4 Loss
TinyGPT-V 2.620 2.343 1.125 1.444
-LoRA 2.620 - Gradient Vanish -
-Input Layer Norm 2.839 2.555 1.344 Gradient Vanish
-RMS Norm 2.747 Gradient Vanish - -
-QK Norm Gradient Vanish - - -
Furthermore, our reveal a notable trend: the smaller the large language model used for transfer learning (particularly in transitioning from text-to-image modality), the more challenging the training process becomes. We observed a pronounced need for additional normalization layers to stabilize the training, especially when scaling down from larger models like Vicuna-13B to smaller ones like Phi-2 (2.7B) and Phi-1.5 (1.3B).

### v2

@@ 4.1 Training
Training process. In our experimental process, we meticulously orchestrated the training of our model through four distinct stages, each characterized by specific learning rate strategies and loss profiles, as shown in Figure 7.
Stage 1: Spanning 17 epochs, with each epoch consisting of 1000 iterations, we employed a dynamic learning rate approach. The learning rate commenced at 1e-5 at the beginning of each epoch and gradually ascended to 1e-4 by the epoch’s end. This pattern was consistently applied across all 17 epochs. The training loss exhibited a steady decline, starting from 7.152 and progressively tapering down to 2.620, reflecting the model’s increasing proficiency in learning from the data. The purpose of this stage is to be able to make the Phi-2 model in TinyGPT-V react in some way to the input of the imaging modality. The alignment of text and image in the semantic space is done.
Stage 2: Comprising 4 epochs, each with 5000 iterations, this stage introduced the “linear_warmup_cosine_lr“ (He et al., 2018; Goyal et al., 2018) learning rate schedule. We initiated a warmup phase of 5000 steps, where the learning rate linearly increased from 1e-6 (warmup_lr) to 1e-4 (init_lr), followed by a cosine decay down to a minimum learning rate of 8e-5. This phase saw a consistent reduction in loss, starting at 2.726 and culminating at 2.343. The purpose of this stage is to enable the LoRA module to play a role in multimodal data, further reducing the model’s loss on image-text pairs and improving the model’s ability to learn from the data.
@@
Method LLM GQA VSR IconVQ VizWiz HM Average
Parameters (zero-shot) (zero-shot) (zero-shot) (zero-shot)
Flamingo 9B - 31.8 - 28.8 57.0 39.20
IDEFICS  (Laurençon et al., 2023) 7B - 38.4 - 35.5 - 37.05
65B - 45.2 - 36.0 - 39.60
BLIP-2 13B 41.0 50.9 40.6 19.6 53.7 41.16
LLaVA 13B 41.3 51.2 43.0 - - 45.17
InstructBLIP 13B 49.5 52.1 44.8 33.4 57.5 47.45
MiniGPT-4 13B - - 35.9 - - 35.90
BLIVA (Hu et al., 2023) 7B - - 44.8 31.4 55.6 41.15
LLaVA-Phi (Zhu et al., 2024) 2.8B - - 54.1 37.6 - 43.15
MoE-LLaVA (Lin et al., 2024) 1.8B×4 61.5 - - 32.6 - 47.50
@@
TinyGPT-V (Phi-2) 2.8B 38.9 54.7 44.7 37.8 54.0 46.02
TinyGPT-V (Phi-1.5) 1.3B 34.3 35.8 37.2 28.4 50.3 37.2
Table 2: Comparative performance of TinyGPT-V and other MLLMs across multiple visual question answering benchmarks. It is worth noting that MoE-LLaVA is required 8xA100-80G for training.
@@ 4.2 Evaluation
Evaluation datasets. GQA (Hudson & Manning, 2019) is a dataset for real-world visual reasoning and compositional question answering, featuring a powerful question engine that generates 22 million diverse reasoning questions. VSR (Liu et al., 2023b) comprises over 10k natural text-image pairs in English, encompassing 66 types of spatial relations. IconQA (Lu et al., 2021) with 107,439 questions aimed at challenging visual understanding and reasoning in the context of icon images, encompassing three sub-tasks (multi-image-choice, multi-text-choice, and filling-in-the-blank). VizWiz (Gurari et al., 2018) is a collection of more than 31,000 visual queries, each derived from a photo taken by a visually impaired individual using a smartphone, accompanied by a vocalized question regarding the image, and supplemented with 10 answers sourced from a crowd for each query. The Hateful Memes dataset (HM) (Kiela et al., 2021), developed by Facebook AI, is a comprehensive multimodal collection specifically designed for the detection of hateful content in memes, combining both image and text elements, and comprises over 10,000 newly created multimodal examples.
TinyGPT-V and others answer example compare
Ures [vqa] where should I hide in this room when playing hide and seek
LLaVA-1.5 hide behind the bookshelf
MiniGPT-v2 behind couch
GPT-4V Behind the Couch Under the Table Inside the Bookshelf Behind the Curtains Behind the TV
TinyGPT-V under couch
Table 3: Comparison of reasoning answers from different Models. Text in red indicates incorrect suggestions. The TinyGPT-V’s answer was short and precise.
@@
Visual question answering results. As shown in Table 2, it becomes evident that TinyGPT-V, a language model with only 2.8 billion parameters, exhibits notably competitive performance across multiple benchmarks, closely rivaling models with nearly 13 billion parameters. Specifically, in the VSR (Visual-Spatial Reasoning) zero-shot task, TinyGPT-V outshines its counterparts by securing the highest score of 54.7%. This is particularly impressive considering its parameter size is approximately 4.6 times smaller than other leading models such as BLIP-2, LLaVA, and InstructBLIP. In the GQA benchmark, while TinyGPT-V scores are 38.9%, it lags behind the highest score achieved by InstructBLIP, which is 49.5%. However, TinyGPT-V shows robust performance in the IconVQ challenge, attaining a score of 44.7%, just 0.1% short of InstructBLIP’s leading score of 44.8%. Similarly, in the VizWiz task, TinyGPT-V demonstrates commendable capabilities with a score of 37.8%, which, is not only the highest but is notable given its reduced parameter count. In the context of the Hateful Memes (HM) dataset, TinyGPT-V matches InstructBLIP’s top score of 57.5% with its own score of 54.0%, again underscoring its efficiency and capacity to compete with models of larger scales. Overall, TinyGPT-V’s performance across these diverse and challenging benchmarks is striking, especially when considering its parameter efficiency
4.3 Qualitative Evaluation
The comparative analysis revealed TinyGPT-V’s distinct advantage in delivering concise and accurate visual interpretations. In the reasoning task to find a hiding spot during a game of hide and seek, TinyGPT-V demonstrated its superior capability by providing a singular, viable suggestion: ’under couch’. This contrasts with other models that either offered multiple options, some of which were incorrect as indicated by the text in red (e.g., GPT-4V suggesting ’Inside the Bookshelf’), or specified less practical hiding spots. When asked about potential activities in an image with an alligator, TinyGPT-V suggested a cautious response without speculating beyond what was visible. In contrast, other models, like LLaVA-1.5, provided extended narratives that introduced assumptions not directly inferred from the image. Similarly, in describing a soccer match scene, TinyGPT-V’s response was succinct and focused on the key elements, avoiding the inaccuracies noted in MiniGPT-v2’s account, which incorrectly identified multiple soccer balls on the pitch. These examples, as tabulated in Table 6 and Table 8, illustrate TinyGPT-V’s superior performance in generating brief yet precise responses, underscoring its practicality for rapid and reliable visual question answering. For efficient evaluation, as shown in table 4, TinyGPT-V operates at the fastest pace, taking only 0.067 seconds to generate a word, which suggests upper efficiency in processing speed compared to LLaVA and MiniGPT-4. On the other hand, LLaVA exhibits a significantly slower word generation time at 0.426 seconds per word, coupled with a higher memory occupancy of 22GB. MiniGPT-4, with a generation time of 0.300 seconds per word and a memory usage of 23.5GB.
Method TinyGPT-V LLaVA MiniGPT-4
seconds per words 0.067 0.426 0.300
inference occupancy (8-bit) 5.6GB 22GB 23.5GB
Table 4: Comparison of inference time and inference occupancy about devices.
@@ 4.4 Ablation Study
Method Stage 1 Loss Stage 2 Loss Stage 3 Loss Stage 4 Loss
TinyGPT-V 2.620 2.343 1.125 1.330
w/o LoRA 2.620 - Gradient Vanish -
w/o Input Layer Norm 2.839 2.555 1.344 Gradient Vanish
w/o RMS Norm 2.747 Gradient Vanish - -
w/o QK Norm Gradient Vanish - - -
@@
Furthermore, our reveal a notable trend: the smaller the large language model used for transfer learning (particularly in transitioning from text-to-image modality), the more challenging the training process becomes. We observed a pronounced need for additional normalization layers to stabilize the training, especially when scaling down from larger models like Vicuna-13B to smaller ones like Phi-2 (2.7B), Phi-1.5 (1.3B), and other small backbones as detailed in the Appendix B.
@@ 5 Conclusion
Appendix A Multi-task Instruction Template
To reduce ambiguity in training a unified multimodal model for diverse tasks such as visual question answering, image captioning, referring expression comprehension, generation, and object parsing and grounding, we employed task-specific tokens from MiniGPT-v2 within a multitask instruction template. This template, derived from the LLaMA-2 conversation template (Touvron et al., 2023), includes a general input format comprising image features, a task identifier token, and an instruction input. We incorporated six distinct task identifiers, each associated with a particular task. For tasks necessitating the identification of spatial locations of referred objects, the model uses textual representations of bounding boxes with coordinates normalized between 0 and 100. Overall, MiniGPT-v2’s unique task-specific tokens enhance task disambiguation, leading to more precise and effective task execution.
TinyGPT-V and others answer example compare
Ures What might happen in this image in the next second
LLaVA-1.5 In the image, a man is sitting on a log next to a large alligator, and he is holding a camera. It is likely that the man is about to take a photograph or record a video of the alligator. The next second could involve the man capturing the alligator’s behavior, appearance, or any other interesting moment. The alligator might also react to the man’s presence, such as moving or showing curiosity, which could make the captured image or video more engaging and memorable.
MiniGPT-v2 In the next second, the man might be startled or surprised by the close-up shot of the crocodile in the water. This could cause him to jump back or react with caution, ensuring his safety in the presence of the wild animal. The image captures a moment of wonder and curiosity, highlighting the beauty of the natural environment while also acknowledging the potential danger that comes with being so close to a wild animal.
GPT-4V In the next second, the man might quickly move away to maintain a safe distance from the alligator. Alligators can be unpredictable, and it’s important to keep a safe distance from them at all times to avoid potential harm. The man seems to be taking a selfie with the alligator, which suggests that he is close to the animal. This is generally not advisable as it can be dangerous. It’s always important to respect wildlife and observe it from a distance that is safe for both the person and the animal.
TinyGPT-V The image shows a man holding a selfie camera next to a crocodile. This could be an unusual and unexpected sight, as crocodiles are not typically associated with human activities like taking selfies. lt’s important for the man to stay safe and avoid any potential risks posed by the crocodile.
Table 6: Comparison of prediction answers from different models. Text in red indicates incorrect suggestions. The TinyGPT-V’s answer was short and precise.
Appendix B Small Backbones for Transfer Learning
As shown in Table 7, a striking pattern emerges from the data: smaller LLMs exhibit heightened sensitivity to the removal of these modules, with a pronounced tendency towards training difficulties, such as gradient vanishing. For instance, the absence of LoRA in both Phi-1.5 and TinyLLaMA resulted in an immediate cessation of training progress post-Stage 1, indicating a critical reliance on this module for sustaining training in smaller models. Similarly, the exclusion of QK Norm led to gradient vanishing at the earliest stage across all smaller LLMs, underscoring its essential role in the initial phases of training. Moreover, the sequential progression in training losses across stages for models without these modifications demonstrates a clear degradation in training efficiency and effectiveness. For example, the removal of Input Layer Norm and RMS Norm not only heightened Stage 1 loss across Phi-1.5 and TinyLLaMA but also precipitated gradient vanishing in later stages, showcasing the compound impact of these modules on model stability and learning capability. This analysis incontrovertibly highlights a fundamental challenge in training smaller LLMs for migration to MLLMs: the absence of key architectural and normalization modules severely impedes their training process, making them more prone to early training halts and efficiency losses. The results underscore the necessity of these components in supporting the stability and gradual learning progression of smaller LLMs, thus illuminating a pivotal consideration for developers aiming to optimize the training framework for seamless model.
LLM Stage 1 Loss Stage 2 Loss Stage 3 Loss Stage 4 Loss
Phi-2 (2.7B) 2.620 2.343 1.125 1.330
Phi-1.5 (1.3B) 3.420 3.043 1.525 1.730
w/o LoRA 3.420 - Gradient Vanish -
w/o Input Layer Norm 3.555 3.221 1.544 Gradient Vanish
w/o RMS Norm 3.557 Gradient Vanish - -
w/o QK Norm Gradient Vanish - - -
TinyLLaMA (1.1B) 3.529 3.053 1.371 1.830
w/o LoRA 3.529 - Gradient Vanish -
w/o Input Layer Norm 3.611 3.331 1.444 Gradient Vanish
w/o RMS Norm 3.607 Gradient Vanish - -
w/o QK Norm Gradient Vanish - - -
Table 7: Importance of each module and different LLMs in TinyGPT-V training framework at each stage of training.
TinyGPT-V and others answer example compare
Ures What’s in the image?
LLaVA-1.5 The image shows a group of men playing soccer on a field. They are actively engaged in the game, with some players running towards the soccer ball.
MiniGPT-v2 The image features a lively and competitive soccer game with several players on the field. Some of the players are wearing jerseys, while others have their own unique outfits. There are at least eight players in total, with three of them being the main focus of the scene. One player is kicking a soccer ball, with another player nearby trying to block the shot. A third player is also present, playing offense. In addition, there are two soccer balls visible on the field, one closer to the left side and the other near the center of the scene. The players are engaged in intense action, showcasing their skills and determination to score. The atmosphere appears to be a high-energy match with a mix of excitement and competition among the players.
GPT-4V The image shows a moment from a soccer match. In the foreground, there’s a player in a blue uniform who appears to be dribbling the ball with close control, evading two opponents in white uniforms. The player in blue is crouched low to the ground, demonstrating agility and focus, which suggests he is in the midst of making a strategic play, possibly attempting to bypass the defenders. In the background, other players can be observed, and the stands are filled with spectators, indicating that this is a professional match likely taking place in a large stadium. The attire and hairstyles suggest the photo may be from a past era, possibly the 1980s or 1990s.
TinyGPT-V The image shows a socer game with several people playing, including at least one person in a blue shirt. The players are engaged in a friendly match, showcasing their skills and teamwork.
Table 8: Comparison of descriptive answers from different Models. Text in red indicates incorrect suggestions. The TinyGPT-V’s answer was short and precise.